OpenStudy (anonymous):

by using the sum and difference identity find the exact value of sin pi/12

OpenStudy (anonymous):

\[\frac{\pi}{3}-\frac{\pi}{4}=\frac{\pi}{12}\] use \[\sin(\alpha -\beta)=\sin(\alpha)\cos(\beta)-\sin(\beta)\cos(\alpha)\] wotj \[\alpha =\frac{\pi}{3},\beta=\frac{\pi}{4}\]

OpenStudy (anonymous):

so what would be the final answer....